OpenStudy (angelab97):

how do you factorize x^4+27x

zepdrix (zepdrix):

Hey Angela! They've both got some x's... factor those out first.\[\large\rm x^4+27x=x(x^3+27)\]From there you want to utilize your `Sum of Cubes Formula`: \(\large\rm a^3+b^3=(a+b)(a^2-ab+b^2)\)

zepdrix (zepdrix):

Hmm but we only have one thing being cubed... x. If you can find a way to rewrite 27 as a perfect cube, you can apply this formula!

OpenStudy (angelab97):

but how do you get (a+b)(a^2-ab+b^2) from x(x^3+27)? @zepdrix

zepdrix (zepdrix):

You get \(\large\rm (a+b)(a^2-ab+b^2)\) from \(\large\rm a^3+b^3\) and you get \(\large\rm a^3+b^3\) from \(\large\rm x^3+27\)

zepdrix (zepdrix):

the x^3 matches up with the a^3, it's in the correct format. But we need to rewrite 27 as \(\large\rm something^3\)

zepdrix (zepdrix):

Example: \(\large\rm 8=2\cdot2\cdot2\) therefore \(\large\rm 8=2^3\) We found a way to write 8 as a perfect cube! Can you find a way to do that with 27?

OpenStudy (angelab97):

thank you!